# Flux-Activated Resonant Control of a Bosonic Quantum Memory

### v1

@@ Abstract
Additionally, we use the multi-tone control to apply arbitrary rotations between two Fock levels. The protocol consists of resonantly transferring the population of a level $|n_{A}+\rangle$ to $|\left(n_{B}-1\right)$-$\rangle$, and then driving a transition from the latter state to $|n_{B}+\rangle$ with an effective rotation angle $\theta$ (Fig. 4a). The sequence is concluded with a second $J_{x}$ applied in reverse, which swaps $|\left(n_{B}-1\right)$-$\rangle$ and $|n_{A}+\rangle$ populations while reverting spectating levels to their original configuration. After the transmon is adiabatically detuned, this operation is equivalent to a Givens rotation between the cavity levels $|n_{A}\rangle$ and $|n_{B}\rangle$ [7]. We show the efficacy of this protocol by applying Givens rotations between Fock state $|1\rangle$ and $|4\rangle$. The parity of the final state is measured as a function of $\theta$, showing a continuous evolution between a minimum value corresponding to $|1\rangle$ and a maximum corresponding to $|4\rangle$ (Fig. 4b). For $\theta=0,\frac{\pi}{4},\frac{\pi}{2}$ and $\pi$, we measure the full Wigner function of the cavity as shown in the plots in Fig. 4c. The density matrix reconstruction of the Wigner measurements indicates state fidelities of $\textnormal{80}\pm\textnormal{2}\%$, $\textnormal{76}\pm\textnormal{2}\%$, $\textnormal{71}\pm\textnormal{2}\%$, $\textnormal{68}\pm\textnormal{2}\%$, respectively. This is consistent with the numerical simulation using real experimental parameter [50].
The performance of the Givens rotation and state preparation are estimated to be limited by transmon dephasing, followed by comparable contributions of $T_{1}$ and coherent errors [50]. The effect of each decoherence channel can be understood in terms of the overlap of JC levels with the corresponding jump operators. The transmon pure dephasing, described by $\sigma_{z}$, translates into a depolarization within each excitation manifold as $\langle N\pm|\sigma_{z}|N\mp\rangle=1$, without introducing pure dephasing as $\langle N\pm|\sigma_{z}|N\pm\rangle=0$. In the limit of the long coherence times of the memory mode, the transmon loss dominates the JC level relaxation. Under this condition, photon loss events are described by $\langle(N-1)\pm|\sigma^{-}|N\pm\rangle=\langle(N-1)\mp|\sigma^{-}|N\pm\rangle=0.5$, leading to an effective JC lifetime twice that of the transmon $T_{1}$, and which does not suffer from bosonic enhancement at higher excitation levels. The modest coherences of this particular transmon also limits the duration of the individual JC drives, which in turn results in coherent leakages due to limited selectivity. A full master equation simulation using these error models is consistent with the experimental observations [50].
@@
There are several known strategies to mitigate the decoherence of flux-tunable transmons that can be adopted to enhance the fidelity of the resonant protocol. For example, the transmon $T_{2}$ can be increased by reducing the flux sensitivity of the SQUID, either by decreasing its area, increasing the asymmetry between the shunting Josephson junctions [20, 50], or using geometries that are inherently more noise-resilient [6]. At the same time, flux modulation can be used to engineer AC sweet spots at the cavity frequency to reduce the direct decoherence limit during the operations. Both strategies are supported by the large bandwidth and strong flux delivery of the on-chip line developed in this work. Besides greater $T_{2}$, increasing the transmon-cavity coupling strength $g$ leads to more selective sideband transitions and faster adiabatic detuning, both of which can reduce the duration of the protocol and mitigate errors. Thus, the performance of the resonant operations shown in this work is not fundamentally limited by the control mechanisms or our device architecture, and can be directly improved by further optimization of device parameters.
@@
In summary, our work demonstrates the control of a long-lived bosonic quantum memory through number-selective transitions in the JC Hamiltonian. This new capability is enabled by a robust flux-tunable architecture that can effectively switch the system between the dispersive and the resonant regimes. We prepare arbitrary superpositions of Fock states via either sequential or multi-tone sideband driving. Moreover, we use efficient state swapping gates to implement arbitrary Givens rotations, providing a path for the universal control of Fock-basis qudits encoded in oscillators [7]. This approach potentially enables the realization of high-dimensional qudits without the increasingly significant charge dispersion that affects higher levels of transmon-based implementations [57]. Furthermore, the resonant control demonstrated here is favorable compared to numerically optimized protocols [15, 14, 12]: in our scheme, the control parameters depend on a finite set of frequencies and amplitudes that can be individually calibrated to high precision without requiring a classical computation step. Additionally, the operations are implemented in a closed subspace, avoiding excursions to higher photon numbers that may amplify loss and nonlinear distortions. Finally, due to the direct access to the JC levels, our protocol leverages a frequency selectivity that scales with the coupling factor $g$, allowing much faster operations compared to other demonstrations of analytical multi-level cavity control [19, 43]. Therefore, our work marks an important contribution to the control capabilities of long-lived microwave cavities in circuit QED and opens up new possibilities for quantum information processing with bosonic modes.
@@
Acknowledgments We acknowledge the funding support from the Singapore Ministry of Education (T2EP50222-0017) and The University of Sydney - National University of Singapore 2026 Ignition Grants. F.V., A.D., N.H., acknowledge the support of the Singapore National Quantum Scholarship Scheme (NQSS). We thank Mr. Kyle Chu and Mr. Jean-Samuel Tettekpoe for their technical inputs during the project development.
Supplementary Material
@@ I Device Design
These flux control features are achieved by combining coplanar waveguide (CPW) and microstrip line architectures. We show the details of the circuit in Fig. 5. The tunable transmon is designed in a microstrip geometry, allowing it to have a large dipole moment that couples strongly with the cavity. The SQUID loop of the transmon is inductively coupled to an on-chip coil biased by a CPW flux line. The flux line is designed with an integrated microwave low-pass filter that mitigates cavity and transmon photon loss.
@@
Figure 5: Chip schematic. It features ancilla DC-SQUID transmon, low-Q readout resonator with Purcell filter, and on-chip flux line with microwave low-pass filter
@@ I.1 Tunable transmon with asymmetric SQUID
Standard transmon [25] qubits feature nonlinear inductance implemented as a Josephson junction, leading to anharmonic oscillator spectrum where one can isolate two-dimensional computational subspace with the transition frequency $\omega_{T}$ given by
$\hbar\omega_{T}=\sqrt{8E_{C}E_{J}}-E_{C},$ (5)
@@
In order to make such a qubit tunable, single Josephson junction should be replaced by a loop with two junctions in parallel, forming a SQUID. Magnetic flux $\Phi_{e}$ threaded through the SQUID loop effectively changes the value of the inductance and, therefore, the qubit frequency, according to
$\hbar\omega_{T}=\sqrt{8E_{C}\cdot 2E_{J}|\cos{\varphi_{e}}|}-E_{C},$ (6)
where $E_{J}$ is the Josephson energy of a single junction, $\varphi_{e}=\pi\frac{\Phi_{e}}{\Phi_{0}}$ the normalised flux , and $\Phi_{0}$ the magnetic flux quantum. Typical spectrum of $\omega_{T}$ as a function of the applied magnetic flux is shown in Fig. 6(a).
Figure 6: Reduced noise sensitivity of asymmetric SQUID transmons. (a) Frequency response for variable junction asymmetry factors $\gamma$. (b) Calculated transmon frequency sensitivity to flux noise as a function of flux bias.
In such tunable devices, noise in the DC bias or in the magnetic environment lead to the fluctuations of the magnetic flux. This, in turn, varies the transmon frequency and leads to significant limitations on the $T_{2}$ lifetime. Thus, it is beneficial to operate at the sweetspots, where the qubit is first-order insensitive to the flux noise. At other bias points, the qubit suffers rapidly increasing dephasing.
This sensitivity to the flux noise can be quantified as $\frac{\partial\omega_{T}}{\partial\Phi_{e}}$, i.e. how much the qubit frequency reacts to fluctuations of the magnetic flux due to the noise. From Eq. (6),
$\frac{\partial\omega_{T}}{\partial\Phi_{e}}=-\frac{\pi^{2}}{\Phi_{0}}\sqrt{2\frac{E_{C}\cdot 2E_{J}}{h^{2}}}\frac{\sin{2\varphi_{e}}}{\cos^{\frac{3}{2}}{\varphi_{e}}}.$ (7)
The sensitivity to flux noise grows rapidly outside of the sweetspots due to the strong frequency dependence on the applied flux bias. Here, we intentionally suppress this dependency by making the two junctions asymmetric, i.e. having different Josephson energies [20]. The extent of asymmetry is expressed as
$\gamma=\frac{E_{J2}}{E_{J1}},$ (8)
@@
Naturally, the spectrum of the asymmetric transmon will also depend on parameter $\gamma$:
$\hbar\omega_{T}=\sqrt{8E_{C}\cdot E_{J1}\sqrt{\gamma^{2}+2\gamma\cos{2\varphi_{e}}+1}}-E_{C}.$ (9)
We calculate the flux noise sensitivity analogously to Eq. (7) using sing (9) and obtain
$\frac{\partial\omega_{T}}{\partial\Phi_{e}}=-\frac{\pi^{2}}{\Phi_{0}}\sqrt{\frac{2E_{C}E_{J1}}{h^{2}}}\frac{4\gamma\sin{2\varphi_{e}}}{\big(\gamma^{2}+2\gamma\cos{2\varphi_{e}}+1\big)^{\frac{3}{4}}}.$ (10)
To compare, we present the simulated flux sensitivies obtained using Eq. (7) and Eq. 10 in Fig. 6(b). One can observe that more asymmetric transmons are less sensitive to the flux noise at the expense of a smaller frequency range. Therefore, for this particular device, we designed the SQUID to be asymmetric with a parameter $\gamma=4.02$. The resulting frequency range is 5.894 - 7.634 GHz, as shown in Fig. 7(a).
It is worth mentioning that Eq. (6) and further analysis, strictly speaking, is only valid for the applied biasing DC current. In case of a AC flux drive, a more rigorous analysis [59, 36] must be applied to ensure an effective differential drive on the SQUID. This optimization can be further complicated by the use of an asymmetric SQUID. although it is outside of the scope of this work, we are actively exploring alternative strategies that can balance the robustness to flux noise with the effective suppression of the common mode drive for future works that require clean AC modulation of the qubit frequency.
@@ I.2 Flux line & on-chip microwave low-pass filter
In this work, we demonstrate the device that combines an on-chip flux line with both effective DC and AC flux delivery with a long-lived cavity. That is achieved by implementing heavy filtering of the flux line: we incorporate compact, third order Chebyshev low-pass filter [41] on the same chip. To ensure effective filtering performance and protection of the cavity mode, we concatenate two such filters, with the stub $Z_{2}$ shared between the two, as depicted schematically in Fig. 5. We leverage the CPW architecture, which allows to reach necessary impedances in much more compact design compared to the microstrip analogue [32], to fit both filtering stages on the 4.1 mm-wide chip. Moreover, our CPW architecture also suppresses cross-talk between the flux and readout modes hosted on the same chip.
The impedance values of the series inductances and stub capacitances used in our filter are listed in Tab. 1, along with the corresponding CPW line dimensions.
@@
Table 1: Impedances of the 3-rd order Chebyshev low-pass filter. Filter diagram is shown in Fig. 2 (a) of the main text.
The simulated response of the on-chip low-pass filter is shown in Fig. 7 (b). The attenuation at the cavity frequency (6.868 GHz) is 98 dB. For the entire transmon frequency band, the attenuation is at least 50 dB.
Figure 7: Filtering and spectrum. (a) A spectrum of the transmon with asymmetric SQUID. Cavity and readout resonators frequency are shown. (b) $S_{21}$ parameter of the cascaded low-pass 3-rd order Chebyshev filter.
@@
Figure 8: Optical images of the device after electron beam lithography. (a) SQUID loop (yellow) with asymmetric junctions (white). (b) Flux line coil, biasing the transmon.
Furthermore, we incorporate a long constriction in the volume where the chip is housed, which we dub the ”attenuation region”. It considerably increases the cutoff frequency of the waveguide structure, enhancing the suppression of the cavity field traveling outside the ground plane. Hence, the cavity field can only reach the drive ports through the CPW mode, which is heavily filtered.

### v2

@@ Abstract
Furthermore, we use the multi-tone control to apply arbitrary rotations between two Fock states. This enacts Givens rotations on a qudit in Fock encoding: $\hat{\mathcal{R}}_{n_{A},n_{B}}(\theta)=\exp[-\imath\theta/2(|n_{A}\rangle\langle n_{B}|+|n_{B}\rangle\langle n_{A}|)]$. Any arbitrary operation on a qudit can be constructed via a combination of $\hat{\mathcal{R}}_{n_{A},n_{B}}(\theta)$ [63, 8], promising a simpler and faster control compared to the previous work [53]. Our implementation consists of resonantly transferring the population of a level $|n_{A}+\rangle$ to $|\left(n_{B}\mathrm{-}1\right)$-$\rangle$, and then driving a transition from the latter state to $|n_{B}+\rangle$ with an effective rotation angle $\theta$ (Fig. 4a). The sequence is concluded with a second $J_{x}$ applied in reverse, which swaps $|\left(n_{B}\mathrm{-}1\right)$-$\rangle$ and $|n_{A}+\rangle$ populations while reverting spectator levels to their original configuration. After the transmon is adiabatically detuned, this operation is equivalent to a Givens rotation between the cavity levels $|n_{A}\rangle$ and $|n_{B}\rangle$ [8]. This operation can be naturally extended to perform gates in any chosen subspace, such as the logical states of the binomial encoding, as we demonstrate in [64]. As a concrete example, we apply Givens rotations $\hat{\mathcal{R}}_{1,4}(\theta)$ between Fock states $|1\rangle$ and $|4\rangle$. The parity of the final state is measured as a function of $\theta$, showing a continuous evolution between a minimum value corresponding to $|1\rangle$ and a maximum corresponding to $|4\rangle$ (Fig. 4b). For $\theta=0,\pi/4,\pi/2$ and $\pi$, we measure the full Wigner function of the cavity as shown in the plots in Fig. 4c. The density matrix reconstruction of the Wigner measurements indicates state fidelities of $\textnormal{80}\pm\textnormal{2}\%$, $\textnormal{76}\pm\textnormal{2}\%$, $\textnormal{71}\pm\textnormal{2}\%$, $\textnormal{68}\pm\textnormal{2}\%$, respectively. This is consistent with the numerical simulation with real experimental parameters [64].
The performance of the Givens rotation and state preparation are limited mainly by transmon dephasing, followed by coherent errors and transmon $T_{1}$ [64]. The effect of each decoherence channel can be understood in terms of the matrix element of JC levels with the corresponding jump operators. Interestingly, transmon pure dephasing does not lead to dephasing of the JC levels, as the corresponding jump operator $\hat{\sigma}_{z}$ has zero matrix element $\langle N\pm|\sigma_{z}|N\pm\rangle=0$. Instead, it translates to depolarisation within each excitation manifold with corresponding matrix element $\langle N\pm|\sigma_{z}|N\mp\rangle=1$.
In the limit of long coherence times of the memory mode, the energy decay of the resonant system is predominantly defined by the transmon relaxation rate corresponding to the jump operator $\hat{\sigma}^{-}$. The matrix elements defining the loss of energy are $\langle(N\mathrm{-}1)\pm|\sigma^{-}|N\pm\rangle=\langle(N\mathrm{-}1)\mp|\sigma^{-}|N\pm\rangle=0.5$, indicating that the energy decay rate of the composite system is half that of the transmon regardless of $N$. In this implementation, the durations of the individual drives are limited by the effective lifetime of the system on resonance. This constraint results in coherent errors caused by leakages to spectator JC levels. A full master equation simulation with these error sources taken into account consistently reproduces experimental observations [64].
@@
To enhance the overall fidelity of our resonant control scheme, we can readily adopt the known strategies to mitigate the decoherence of flux-tunable transmons, which is the dominant source of imperfection in this implemenation. For example, the transmon $T_{2}$ can be increased by reducing the flux sensitivity of the SQUID either by decreasing its area, increasing the asymmetry between the shunting Josephson junctions [24, 64], or using geometries that are inherently more noise-resilient [7]. At the same time, flux modulation can be used to engineer AC sweet spots at the cavity frequency to reduce the direct decoherence limit during the operations. Both strategies are supported by the large bandwidth and strong flux delivery of the on-chip line developed in this work. Besides greater $T_{2}$, increasing the transmon-cavity coupling strength $g$ leads to more selective sideband transitions and faster adiabatic detuning, both of which can reduce the duration of the protocol and mitigate errors. Thus, the performance of the resonant operations shown in this work is not fundamentally limited by the control mechanisms or our device architecture, and can be directly improved by further optimisation of device parameters.
@@
In summary, our work demonstrates the control of a long-lived bosonic quantum memory through number-selective transitions in the JC Hamiltonian. This new capability is enabled by a robust flux-tunable architecture that can effectively switch the system between the dispersive and the resonant regimes. We prepare arbitrary superpositions of Fock states via either sequential or multi-tone sideband driving. Moreover, we use the JC transitions to efficiently implement arbitrary Givens rotations, providing a path for the universal control of qudits encoded in oscillators. Finally, thanks to the direct access to the JC levels, our protocol is intuitive and calibration-friendly. Moreover, it leverages a frequency selectivity that scales with the coupling factor $g$, allowing potentially faster control compared to typical dispersive operations and other parametrically activated techniques. Therefore, our work marks an important step towards more efficient control capabilities of long-lived microwave quantum memories in cQED and opens up new possibilities for quantum information processing with bosonic modes.
@@
Acknowledgements We acknowledge the funding support from the Singapore Ministry of Education (T2EP50222-0017) and The University of Sydney - National University of Singapore 2026 Ignition Grants. F.V., A.D., N.H., acknowledge the support of the Singapore National Quantum Scholarship Scheme (NQSS). We thank Mr. Kyle Chu and Mr. Jean-Samuel Tettekpoe for their technical inputs during the project development. We thank Dr. Yao Lu and Dr. Mustafa Bakr for the valuable discussions during the writing of this manuscript.
Supplementary Materials
Flux-Activated Resonant Control of a Bosonic Quantum Memory
@@ I Device Design
These flux control features are achieved by combining coplanar waveguide (CPW) and microstrip line architectures. We show the details of the circuit in Fig. S1. The tunable transmon is designed in a microstrip geometry, allowing it to have a large dipole moment that couples strongly to the cavity. The SQUID loop of the transmon is inductively coupled to an on-chip coil biased by a CPW flux line. The flux line is designed with an integrated microwave low-pass filter that mitigates cavity and transmon photon loss.
@@
Figure S1: Chip schematic. It features ancilla DC-SQUID transmon, low-Q readout resonator with Purcell filter, and on-chip flux line with microwave low-pass filter
@@ I.1 Tunable transmon with asymmetric SQUID
Standard transmon [31] qubits feature nonlinear inductance implemented as a Josephson junction, leading to anharmonic oscillator spectrum where one can isolate two-dimensional computational subspace with the transition frequency $\omega_{\textnormal{T}}$ given by
$\hbar\omega_{\textnormal{T}}=\sqrt{8E_{C}E_{J}}-E_{C},$ (S1)
@@
In order to make such a qubit tunable, single Josephson junction should be replaced by a loop with two junctions in parallel, forming a DC-SQUID. Magnetic flux $\Phi_{e}$ threaded through the SQUID loop effectively changes the value of the inductance and, therefore, the qubit frequency, according to
$\hbar\omega_{\textnormal{T}}=\sqrt{8E_{C}\cdot 2E_{J}|\cos{\varphi_{e}}|}-E_{C},$ (S2)
where $E_{J}$ is the Josephson energy of a single junction, $\varphi_{e}=\pi\frac{\Phi_{e}}{\Phi_{0}}$ the normalised flux , and $\Phi_{0}$ the magnetic flux quantum. Typical spectrum of $\omega_{\textnormal{T}}$ as a function of the applied magnetic flux is shown in Fig. S2(a).
Figure S2: Reduced noise sensitivity of asymmetric SQUID transmons. (a) Frequency response for variable junction asymmetry factors $\gamma$. (b) Calculated transmon frequency sensitivity to flux noise as a function of flux bias.
In such tunable devices, noise in the DC bias or in the magnetic environment leads to the fluctuations of the magnetic flux. This, in turn, varies the transmon frequency and leads to significant limitations on the $T_{2}$ lifetime. Thus, it is beneficial to operate at the sweetspots, where the qubit is first-order insensitive to the flux noise. At other bias points, the qubit suffers rapidly increasing dephasing.
This sensitivity to the flux noise can be quantified as $\frac{\partial\omega_{\textnormal{T}}}{\partial\Phi_{e}}$, i.e. how much the qubit frequency reacts to the fluctuations of the magnetic flux due to the noise. From Eq. (S2),
$\frac{\partial\omega_{\textnormal{T}}}{\partial\Phi_{e}}=-\frac{\pi^{2}}{\Phi_{0}}\sqrt{2\frac{E_{C}\cdot 2E_{J}}{h^{2}}}\frac{\sin{2\varphi_{e}}}{\cos^{\frac{3}{2}}{\varphi_{e}}}.$ (S3)
The sensitivity to the flux noise grows rapidly outside of the sweetspots due to the strong dependence of frequency on the applied flux bias. Here, we intentionally suppress this dependency by making the two junctions asymmetric, i.e. having different Josephson energies [25]. The extent of asymmetry is expressed as
$\gamma=\frac{E_{J2}}{E_{J1}},$ (S4)
@@
Naturally, the spectrum of the asymmetric transmon will also depend on the parameter $\gamma$:
$\hbar\omega_{\textnormal{T}}=\sqrt{8E_{C}\cdot E_{J1}\sqrt{\gamma^{2}+2\gamma\cos{2\varphi_{e}}+1}}-E_{C}.$ (S5)
We calculate the flux noise sensitivity analogously to Eq. (S3) using Eq. (S5) and obtain
$\frac{\partial\omega_{\textnormal{T}}}{\partial\Phi_{e}}=-\frac{\pi^{2}}{\Phi_{0}}\sqrt{\frac{2E_{C}E_{J1}}{h^{2}}}\frac{4\gamma\sin{2\varphi_{e}}}{\big(\gamma^{2}+2\gamma\cos{2\varphi_{e}}+1\big)^{\frac{3}{4}}}.$ (S6)
We present the simulated flux sensitivies obtained using Eq. (S3) and Eq. (S6) in Fig. S2(b). One can observe that more asymmetric transmons are less sensitive to the flux noise at the expense of a smaller frequency range. Therefore, for this particular device, we designed the SQUID to be asymmetric with a parameter $\gamma=4.02$. The resulting frequency range is 5.894 - 7.634 GHz, as shown in Fig. S3(a).
It is worth mentioning that Eq. (S2) and further analysis, strictly speaking, is only valid for the applied biasing DC current. In case of an AC flux drive, a more rigorous analysis [74, 44] must be applied to ensure an effective differential drive on the SQUID. The use of an asymmetric SQUID makes this optimisation more complicated. Although it is outside of the scope of this work, we are actively exploring alternative strategies that can balance the robustness to flux noise with the effective suppression of the common mode drive for future works that require clean AC modulation of the qubit frequency.
@@ I.2 Flux line & on-chip microwave low-pass filter
In this work, we demonstrate the device that combines an on-chip flux line with both effective DC and AC flux delivery with a long-lived cavity. That is achieved by implementing heavy filtering of the flux line: we incorporate compact, third order Chebyshev low-pass filter [56] with cutoff frequency of 4 GHz on the same chip. To ensure effective filtering performance and protection of the cavity mode, we concatenate two such filters, with the stub $Z_{2}$ shared between the two, as depicted schematically in Fig. S1. We leverage the CPW architecture, which allows to reach necessary impedances in much more compact design compared to the microstrip analogue [40] and to fit both filtering stages on the 4.1 mm-wide chip. Moreover, our CPW architecture also suppresses cross-talk between the flux and readout modes hosted on the same chip.
The impedance values of the series inductances and stub capacitances used in our filter are listed in Tab. I, along with the corresponding CPW line dimensions.
@@
Table I: Impedances of the 3-rd order Chebyshev low-pass filter. Filter diagram is shown in Fig. 2 (a) of the main text.
The simulated response of the on-chip low-pass filter is shown in Fig. S3 (b). The attenuation at the cavity frequency (6.868 GHz) is 98 dB. For the entire transmon frequency band, the attenuation is at least 50 dB.
Figure S3: Filtering and spectrum. (a) A spectrum of the transmon with asymmetric SQUID. Cavity and readout resonators frequency are shown. (b) $S_{21}$ parameter of the cascaded low-pass 3-rd order Chebyshev filter.
@@
Figure S4: Optical images of the device after electron beam lithography. (a) SQUID loop (yellow) with asymmetric junctions (white). (b) Flux line coil, biasing the transmon.
Furthermore, we incorporate a long constriction in the volume where the chip is housed, which we dub the "attenuation region". It considerably increases the cutoff frequency of the waveguide structure, enhancing the suppression of the cavity field traveling outside the ground plane. Hence, the cavity field can only propagate to the drive ports through the CPW mode, which is heavily filtered.